OpenStudy (vincentblackshadow):

The total price (P) to order a certain number of team jerseys (J) is represented by the following equation:        P = 20J 11) Which of the following accurately represent the equation?

OpenStudy (vincentblackshadow):

A) Price Jerseys 20 400 30 600 40 800 B) Jerseys Price 100 1000 200 2000 400 400 C) Jerseys Price 20 400 30 600 40 800

OpenStudy (vincentblackshadow):

D) None of these graphs represent the equation

OpenStudy (jordanboss):

should be B

OpenStudy (vincentblackshadow):

thank you, would you mind helping me with a few more?

OpenStudy (jordanboss):

ok

OpenStudy (vincentblackshadow):

this is from the same question above

OpenStudy (vincentblackshadow):

) Which of the following accurately represent the equation? A) f(J) = P B) Price is a function of number of jerseys . C) Number of jerseys is a function of price. D) (40, 1600),  (20, 800)  (10, 200) E) f(P) = J

OpenStudy (jordanboss):

C

OpenStudy (vincentblackshadow):

Find the slope. 21) **  y = 10 3 x + 8

OpenStudy (vincentblackshadow):

hold on

OpenStudy (vincentblackshadow):

i typed that wrong

OpenStudy (vincentblackshadow):

its y = 10/3x+ 8

OpenStudy (jordanboss):

y=2

OpenStudy (vincentblackshadow):

Find the coordinates of the y-intercept. 10 22) y = 9 x + 7 Find the coordinates of the x-intercept. 23) x + y = -5 Graph the linear equations. 24) y = 3x + 1 25) 2x - y = -8

OpenStudy (vincentblackshadow):

disregaurd number 24

OpenStudy (jordanboss):

ok hold on